# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

### v3

@@ Abstract.
Data missingness is a practical challenge of sustained interest to the scientific community. In this paper, we present Shades-of-Null, an evaluation suite for responsible missing value imputation. Our work is novel in two ways (i) we model realistic and socially-salient missingness scenarios that go beyond Rubin’s classic Missing Completely at Random (MCAR), Missing At Random (MAR) and Missing Not At Random (MNAR) settings, to include multi-mechanism missingness (when different missingness patterns co-exist in the data) and missingness shift (when the missingness mechanism changes between training and test) (ii) we evaluate imputers holistically, based on imputation quality, as well as on the predictive performance, fairness and stability of the models that are trained and tested on the data post-imputation.
We use Shades-of-Null to conduct a large-scale empirical study involving 23,940 experimental pipelines, and find that while there is no single best-performing imputation approach for all missingness types, interesting trade-offs arise between predictive performance, fairness and stability, based on the combination of missingness scenario, imputer choice, and the architecture of the predictive model. We make Shades-of-Null publicly available, to enable researchers to rigorously evaluate missing value imputation methods on a wide range of metrics in plausible and socially meaningful scenarios.
@@ 1. Introduction
Specifically in data management, debates have been raging since the very inception of the field regarding whether — and how — to handle missing values, see  Date (1984) for a classic example. One difficulty, at the operational level, is that, while missing values are typically denoted with the token null, that may not always be the case. There could be hidden missing values in the data, like when ‘AL’ is selected as the state on a job application because it’s the first item in the drop-down list. Another difficulty, at the semantic level, is that a null can take on many meanings: it may signify that the value is unknown, inapplicable, intentionally withheld, etc.
In this paper, we do not concern ourselves with this semantic debate, and we also do not consider hidden missing values (Qahtan et al., 2018). Rather, we focus our attention on a very specific case. We are given a dataset $X$ (a single relation), in which the values of some features are missing, denoted by the marker null. The meaning of null is that the feature does take on a value in the real world (i.e., it is applicable) but that value is unobserved in $X$. Further, we intend to use $X$ in a machine learning (ML) setting: to train a model or to make predictions about the tuples in $X$ using a model that has already been trained and deployed. This model cannot deal with null values directly, and so these values must be imputed (filled in) as part of data pre-processing.
@@
Consider a dataset of job applicants that includes their most recent salary and the number of years of relevant work experience. The MCAR assumption holds, for example, when a job applicant’s most recent salary is missing from their record for purely administrative reasons that are uncorrelated with either their (potentially missing) salary or with their (observed) experience. More formally, the probability that a feature value is missing is independent of $X$: $P(R|X)=P(R)$.
@@
Suppose that job applicants with fewer years of relevant work experience are more likely to withhold their most recent salary (in hopes of being offered a higher salary by the prospective employer), and that this can be explained by observed covariates (i.e., by the number of years of work experience). The MAR assumption holds in this example, where the probability that some feature value is missing depends only on the observed covariates $X_{\text{obs}}$, and not on the missing feature values: $P(R|X)=P(R|X_{\text{obs}})$.
@@
Suppose that a job applicant’s most recent salary does not correlate with their number of years of experience but rather is based on some combination of their geographic location and the results of a skills test, neither of which are captured in the data. Further, suppose that individuals whose most recent salary was low may be likely to withhold their salary information (again, in hopes of being offered a higher salary by the prospective employer). The MNAR assumption holds in this example, because missingness is correlated with the missing value itself, and it cannot be explained by the observed covariates (i.e., by the number of years of experience): $P(R|X)\not=P(R|X_{\text{obs}})$.
@@
Rubin’s framework has formed the basis of a large body of technical work on missing value imputation, which has been covered in several excellent and comprehensive surveys (Abdelaal et al., 2023; Hasan et al., 2021; Le Quy et al., 2022; Mena et al., 2022; Pintz et al., 2022; Emmanuel et al., 2021; Peherstorfer et al., 2018; Lin and Tsai, 2020; Pigott, 2010; Zhou et al., 2024; Alabadla et al., 2022; García Laencina et al., 2010; Joel and Doorsamy, 2022; Ilyas and Chu, 2019). Broadly speaking, MVI approaches fall into several categories: (1) statistical, such as imputing missing observations with the median or mode (Schafer and Graham, 2002); (2) learning-based impute-then-classify, which use the observed data to iteratively learn and impute missing values, using k-nearest neighbors (Batista and Monard, 2003), clustering (Gajawada and Toshniwal, 2012), decision trees (Twala, 2009), or ensembles (Stekhoven and Bühlmann, 2012); and (3) joint data cleaning and model training (Karlaš et al., 2020; Krishnan et al., 2017, 2016), based on Rubin’s seminal multiple imputation framework (Rubin, 1987).
@@
Rubin’s framework, while clean and amenable to analysis, does not fully capture real-world missingness. First, it has been argued that the assumptions behind MCAR rarely hold in practice, and that most real-world data exists on a continuum between MAR and MNAR, depending on how that data was collected (Graham, 2009). Second, while it is useful to distinguish between missingness scenarios based on their statistical properties to draw valid inferences, we are likely to encounter multi-mechanism missingness in practice, where several scenarios may co-exist in the same dataset, affecting subsets of its features, tuples, or both (Zhou et al., 2024). For example,  Mitra et al. (2023) introduce the notion of a data missingness life cycle and posit that data pre-processing steps such as assimilating data from different sources or from different heterogeneous populations (each with their own missingness patterns) can result in ‘structured missingness’ that cannot be represented using Rubin’s framework. Third, it has been observed that, in the context of data-centric AI, the assumptions that hold about missingness in the training data may no longer hold post-deployment, a phenomenon termed missingness shift (by analogy with data distribution shift) (Zhou et al., 2023).
Evaluating MVI techniques.
Despite an abundance of MVI techniques using different model architectures being proposed each year, making meaningful progress on learning from incomplete data ultimately comes down to how realistically we model missingness: our proposed methods are only as good as the evaluation scenarios we rigorously test them on. If our evaluation scenarios do not realistically capture the kinds of missingness that we encounter in practice, then our data cleaning methods are unlikely to be effective. Further, newly proposed MVI techniques are often only evaluated on the correctness of the imputation (imputation quality) and the correctness of the predictive model (accuracy or F1). However, in recent years, the scientific community is starting to uncover trade-offs between MVI efficacy and other normatively important dimensions of model performance, including algorithmic fairness (Guha et al., 2024; Feng, 2023), and stability and robustness (Zhou et al., 2024), beyond accuracy.
Missingness as a form of pre-existing bias.
@@
Consider again the job applicant screening example, and suppose that the applicants’ gender and age are among the features. Female job applicants who suspect that they have been receiving lower salaries than their male colleagues, despite having comparable qualifications and job experience, may withhold salary information more frequently then men, hoping to narrow the gender wage gap in their next position. As a result there will be more missing salary values for women, and the probability that the value is missing depends on the observed covariate, gender. This situation is consistent with MAR missingness, and it points to pre-existing bias — a situation where data reflects a history of societal discrimination (Friedman and Nissenbaum, 1996). For our final example, suppose that disability status is among the features. Job applicants who live with a disability are more likely to omit a value for this feature. Assuming that this feature does not correlate with the other features, this situation is consistent with MNAR, with missingness itself serving as a proxy for disadvantage.
@@
Our work is (1) novel: to the best of our knowledge, the settings of multi-mechanism missingness and missingness shifts have not been empirically studied before; (2) comprehensive: we evaluate a suite of 8 MVI techniques on 7 benchmark datasets using 6 model types, running a total of 23,940 pipelines, and is the first study of such scale in the missing data domain, to the best of our knowledge; (3) normatively grounded: we focus on decision-making contexts such as lending, hiring, and healthcare, where missingness is socially salient. Mitigating social harm such as algorithmic discrimination is a leading concern in these domains (Barocas and Selbst, 2016), and we evaluate the impact of MVI approaches on downstream model fairness and stability (which have been understudied in the context of missing data), in addition to classically studied imputation quality and model correctness metrics.
While developing the Shades-of-Null evaluation suite, we found and fixed several bugs in existing MVI implementations, including data leakage and omitted hyperparameter tuning. See Appendix A.2 for details.
We make Shades-of-Null publicly available [^1] https://github.com/FalaahArifKhan/data-cleaning-stability and hope to enable researchers to comprehensively and rigorously evaluate new MVI methods on a wide range of evaluation metrics, under plausible and socially meaningful missingness scenarios.
@@ 2. Related Work
Learning-based approaches have become increasingly popular, and include k-nearest neighbors, decision trees, support vector machines, clustering, and ensembles (Batista and Monard, 2003; Lin and Tsai, 2020; Hasan et al., 2021).  Zhou et al. (2024) and Liu et al. (2023) review deep learning-based approaches (variational auto-encoders and generative adversarial networks) and representation learning (graph neural networks and diffusion-based methods). Multiple imputation (Schafer and Graham, 2002; Zhou et al., 2024) and expectation maximization (Pigott, 2010; Wang et al., 2024) are also influential, but too computationally expensive to be popular in practice (Alabadla et al., 2022).
@@
Empirical studies have been overwhelmingly focused on medical datasets, and only evaluate missingness under MCAR (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008). Further, most proposed methods only perform direct evaluation, using metrics such as MAE, MSE, RMSE, and AUC (Hasan et al., 2021; Joel and Doorsamy, 2022). Very few studies use both direct and indirect evaluation, and even studies that include indirect evaluation only evaluate predictive accuracy (Lin and Tsai, 2020). This leaves a significant research gap, specially for desiderata such as stability and self-consistency, which have recently been shown to be highly sensitive to data-centric interventions (Black and Fredrikson, 2021; Black et al., 2022; Cooper et al., 2023), and unfairness, which can be worsened by improperly chosen data cleaning techniques (Guha et al., 2024). Additionally, the performance of MVI techniques under multi-mechanism missingness (Zhou et al., 2024) and missingness shifts (Zhou et al., 2023) remains unexplored, despite these conditions being more likely to occur in practice due to distribution shifts in production deployments (Graham, 2009).
Notably, and worryingly, overwhelming evidence in the literature indicates that there is no single “best-performing” MVI approach on accuracy  (Hasan et al., 2021; Lin and Tsai, 2020; Shadbahr et al., 2023; García Laencina et al., 2010; Farhangfar et al., 2008), and that model performance (narrowly measured based on ‘correctness’ thus far) depends on dataset characteristics such as size and correlation between variables (Alabadla et al., 2022) and missingness rates in the train and test sets (Lin and Tsai, 2020; Shadbahr et al., 2023).
@@
There has been some recent interest in studying the social harm that can come from poorly chosen MVI techniques (Caton et al., 2022; Zhang and Long, 2022, 2021a, 2021b; Jeong et al., 2022; Wang and Singh, 2021; Martínez-Plumed et al., 2019; Feng et al., 2024). Most empirical studies (Wang and Singh, 2021; Zhang and Long, 2021b, 2022; Caton et al., 2022) have worked with the COMPAS (Larson et al., 2016) and Adult  [^2] https://archive.ics.uci.edu/dataset/2/adult datasets, the latter of which has been “retired” from community use due to issues with provenance (Ding et al., 2021). Further, these experimental studies employ randomly-generated missingness: usually by randomly sampling or using a fixed set of columns, and randomly picking rows in which to replace values with null. We critique this approach, since detecting and mitigating unfairness requires broader socio-technical thinking, such as having higher rates of missingness for minority groups and in features that are highly correlated with sensitive attributes (called proxy variables in the fairness literature) (Caton et al., 2022).
@@
A notable exception is  Martínez-Plumed et al. (2019), who map social mechanisms such as prejudicial access and self-reporting bias to different categories of missingness such as missing-by-design and item non-response. Further, they analyse the correlations between all features to study the effect of different kinds of missingness. We use a similar methdology to simulate realistic missingnesses in this work. Importantly, however, we find conceptual limitations in their framing of algorithmic fairness. Martínez-Plumed et al. (2019) write:“The surprising result was to find that, […] the examples with missing values seem to be fairer than the rest.” Asserting that some rows of data are more or less fair does not make sense, in that samples (e.g., job applicants) do not demonstrate the property of being fair. It is the model (e.g., used to make hiring decisions) that asserts the property of being fair or not, by the inclusion or exclusion of certain samples amongst those given a positive outcome. We re-interpret the findings of  Martínez-Plumed et al. (2019) to say that excluding samples with missingness can increase model unfairness, which is further evidence against the use of deletion as a strategy for dealing with missing data.
@@
Zhang and Long (2021b) evaluate MVI methods on imputation fairness, defined as the imputation accuracy difference on privileged and disadvantaged demographic groups. They find that (1) higher missingness disparity leads to higher imputation unfairness; (2) higher rates of missingness lead to higher imputation unfairness; (3) more imbalance in the data (or larger disparity in the sample sizes across groups) leads to higher imputation unfairness; (4) the same imputation method has different prediction fairness under different missingness mechanisms. Their evaluation is limited to randomly generated null values in COMPAS. We extend their analysis of imputation fairness to additional datasets, missingness scenarios and imputation fairness definitions (such as KL divergence difference).
@@
Zhang and Long (2022) introduce the notion of imputation fairness risk and provide bounds for “correctly specified” imputation methods. While there is very little theoretical work in this space, and this is a commendable contribution, we call into question the broader implications of this result: it is well-known that imputation quality metrics do not fully capture downstream model performance (Zhang and Long, 2021a). Put differently, a classifier built on poor imputation quality can still give good downstream performance (Shadbahr et al., 2023). With this in mind, we ask the question: Does minimizing imputation unfairness minimize model unfairness? Our empirical findings indicate it does not, as is discussed in Section 4.4.
@@
Missingness and stability. We are not aware of any work investigating the effect of missing value imputation on model stability.

### v4

@@ Abstract.
Data missingness is a practical challenge of sustained interest to the scientific community. In this paper, we present Shades-of-Null, an evaluation suite for responsible missing value imputation. Our work is novel in two ways (i) we model realistic and socially-salient missingness scenarios that go beyond Rubin’s classic Missing Completely at Random (MCAR), Missing At Random (MAR) and Missing Not At Random (MNAR) settings, to include multi-mechanism missingness (when different missingness patterns co-exist in the data) and missingness shift (when the missingness mechanism changes between training and test) (ii) we evaluate imputers holistically, based on imputation quality and imputation fairness, as well as on the predictive performance, fairness and stability of the models that are trained and tested on the data post-imputation.
We use Shades-of-Null to conduct a large-scale empirical study involving 29,736 experimental pipelines, and find that while there is no single best-performing imputation approach for all missingness types, interesting trade-offs arise between predictive performance, fairness and stability, based on the combination of missingness scenario, imputer choice, and the architecture of the predictive model. We make Shades-of-Null publicly available, to enable researchers to rigorously evaluate missing value imputation methods on a wide range of metrics in plausible and socially meaningful scenarios.
@@ 1. Introduction
Debates on handling missing values in data management date back to the field’s inception, with classic discussions such as Date (1984). At the operational level, missing values are typically denoted by null, but hidden missing values can exist (e.g., ‘AL’ being selected by default in a job application). At the semantic level, null can have multiple meanings—unknown, inapplicable, or intentionally withheld. This paper does not engage in the semantic debate or consider hidden missing values (Qahtan et al., 2018). Instead, we focus on a specific case: a dataset $X$ (a single relation) where some features are missing, marked by null, indicating that the feature has a real-world value but is unobserved in $X$. Our goal is to use $X$ in a machine learning (ML) setting, either for model training or inference. Since ML models cannot handle null directly, missing values must be imputed as part of data preprocessing.
@@
In a job applicant dataset with salary and years of experience, MCAR holds if salary is missing due to administrative errors, unrelated to the salary itself or work experience. That is: $P(R|X)=P(R)$.
@@
If job applicants with fewer years of experience are more likely to withhold their salary, and this can be explained by observed covariates (i.e., years of experience), then MAR holds. Here, missingness depends only on observed features, not the missing values themselves: $P(R|X)=P(R|X_{\text{obs}})$.
@@
Consider a job applicant whose salary depends on geographic location and skills test results—neither captured in the data—rather than years of experience. Suppose applicants with lower salaries are more likely to withhold this information, hoping for a higher offer. In this case, MNAR holds because missingness is correlated with the missing value itself and cannot be explained by observed covariates (i.e., years of experience): $P(R|X)\not=P(R|X_{\text{obs}})$.
@@
Rubin’s framework has shaped a vast body of work on missing value imputation, extensively reviewed in several comprehensive surveys (Abdelaal et al., 2023; Hasan et al., 2021; Le Quy et al., 2022; Mena et al., 2022; Pintz et al., 2022; Emmanuel et al., 2021; Peherstorfer et al., 2018; Lin and Tsai, 2020; Pigott, 2010; Zhou et al., 2024; Alabadla et al., 2022; García Laencina et al., 2010; Joel and Doorsamy, 2022; Ilyas and Chu, 2019). MVI methods fall into three main categories: (1) Statistical methods, such as median or mode imputation (Schafer and Graham, 2002); (2) Learning-based impute-then-classify, which iteratively impute missing values using k-nearest neighbors (Batista and Monard, 2003), clustering (Gajawada and Toshniwal, 2012), decision trees (Twala, 2009), or ensembles (Stekhoven and Bühlmann, 2012); (3) Joint data cleaning and model training, integrating imputation with model learning (Karlaš et al., 2020; Krishnan et al., 2017, 2016), based on Rubin’s multiple imputation framework (Rubin, 1987).
@@
Rubin’s framework, while analytically clean, does not fully capture real-world missingness. First, MCAR assumptions rarely hold, and real-world data often falls on a continuum between MAR and MNAR, depending on collection methods (Graham, 2009). Second, missingness mechanisms frequently co-exist within a dataset (affecting different features or tuples), leading to multi-mechanism missingness (Zhou et al., 2024). For instance, Mitra et al. (2023) introduce the data missingness life cycle, showing how data integration from diverse sources creates structured missingness beyond Rubin’s model. Third, in data-centric AI, missingness assumptions valid during training may shift post-deployment, a phenomenon termed missingness shift, analogous to data distribution shift (Zhou et al., 2023).
Missingness as a form of bias.
@@
Consider the job applicant screening example with gender and age as features. Female applicants who suspect wage discrimination may withhold salary information more often than men, hoping to narrow the gender pay gap. This leads to more missing salary values for women, where missingness depends on the observed covariate (gender), aligning with MAR. This reflects pre-existing bias, where data encodes historical societal discrimination (Friedman and Nissenbaum, 1996). For another example, suppose disability status is included as a feature. Applicants with disabilities may be more likely to omit this information. If disability status is uncorrelated with other features, this scenario aligns with MNAR, with missingness itself acting as a proxy for disadvantage.
When handling missing values, data scientists must also address technical bias (Friedman and Nissenbaum, 1996), where incorrect technical choices create disparities in predictive accuracy, often amplifying pre-existing bias. A key example is imputing missing values under incorrect assumptions, which can worsen disparities in classifier performance (Schelter and Stoyanovich, 2020; Stoyanovich et al., 2022; Schelter et al., 2019; Guha et al., 2024). For instance, if job applicant salaries are missing under MAR or MNAR (e.g., older women withhold salaries due to perceived discrimination), imputing them under MCAR could further depress salary estimates, reinforcing the gender wage gap and ageism, and leading to discriminatory outcomes.
Missing value imputation can impact model arbitrariness.
Missingness is an indication of uncertainty in the data. MVI methods “resolve” this uncertainty at the tuple level, but they may induce a change in the data distribution in ways that impacts the stability of predictions of a model trained on this data. In some cases, the resulting models produce vastly different — and even arbitrary — predictions under small perturbations in the input (Coston et al., 2023; Creel and Hellman, 2022; Rhea et al., 2022b, a). For example, if a job applicant’s salary is imputed in vastly different ways upon two consecutive applications for the same position, and this, in turn, impacts the hiring decision, then the decision-making process violates the principle of process fairness (e.g., (sta, 1983; Tyler, 2006)). Importantly, instability and accuracy are orthogonal: models can be accurate in expectation while still being unstable (Long et al., 2023).
Research gap
Despite numerous MVI techniques being proposed each year, there has been limited systematic progress in assessing them across key performance aspects, including imputation correctness, predictive accuracy, and fairness—measured as disparities in imputation quality or model performance across groups. Moreover, while missingness signals uncertainty, there has been no comprehensive evaluation of the stability of models trained on cleaned data. Crucially, realistic modeling of missingness, identifying bias sources, and selecting appropriate stakeholder groups and fairness metrics must be grounded in the specific context of use (Mitchell et al., 2021; Fuster et al., 2022; Obermeyer et al., 2019; Köchling and Wehner, 2020). For instance, age-based discrimination is relevant in both hiring and lending, yet older applicants face disadvantages (and legal protections) in hiring, while younger applicants are disadvantaged in lending. Thus, MVI techniques must be evaluated in societally meaningful scenarios.
@@
Our work is (1) novel: to the best of our knowledge, the settings of multi-mechanism missingness and missingness shifts have not been empirically studied before; (2) comprehensive: we evaluate a suite of 15 MVI techniques on 7 benchmark datasets using 6 model types, running a total of 29,736 pipelines, and is the first study of such scale in the missing data domain, to the best of our knowledge; (3) normatively grounded: we focus on decision-making contexts such as lending, hiring, and healthcare, where missingness is socially salient. Mitigating social harm such as algorithmic discrimination is a leading concern in these domains (Barocas and Selbst, 2016), and we evaluate the impact of MVI approaches on downstream model fairness and stability (which have been understudied in the context of missing data), in addition to classically studied imputation quality and model correctness metrics.
While developing the Shades-of-Null evaluation suite, we found and fixed several bugs in existing MVI implementations, including data leakage and omitted hyperparameter tuning. See Appendix A.2 for details. We make Shades-of-Null publicly available  [^1] https://github.com/FalaahArifKhan/data-cleaning-stability and hope to enable researchers to comprehensively evaluate new MVI methods on a wide range of evaluation metrics, under plausible and socially meaningful missingness scenarios.
@@ 2. Related Work
Learning-based approaches have become increasingly popular, and include k-nearest neighbors, decision trees, support vector machines, clustering, and ensembles (Batista and Monard, 2003; Lin and Tsai, 2020; Hasan et al., 2021).  Zhou et al. (2024) and Liu et al. (2023) review deep learning-based approaches (variational auto-encoders and generative adversarial networks) and representation learning (graph neural networks and diffusion-based methods). Multiple imputation (Schafer and Graham, 2002; Zhou et al., 2024) and expectation maximization (Pigott, 2010; Wang et al., 2024a) are also influential, but too computationally expensive to be popular in practice (Alabadla et al., 2022).
MNAR-specific techniques, like not-MIWAE (Ipsen et al., 2020) and GINA (Ma and Zhang, 2021), tackle the challenge of MNAR data by employing identifiable generative models that effectively account for complex missingness mechanisms. Recent methods, including NOMI (Wang et al., 2024b) and TDM (Zhao et al., 2023), introduce advancements like uncertainty-driven networks and transformed distribution matching, which enhance both imputation accuracy and computational efficiency.
@@
We are aware of several surveys of MVI techniques, all conducted with a strong empirical focus (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008; Miao et al., 2022; Shahbazian and Greco, 2023). Miao et al. (2022) compare 19 MVI methods on 15 datasets, and while our results corroborate their findings (see Section 5.1), their evaluation is limited to imputation quality and overall accuracy (but not fairness or stability). Other empirical studies have been primarily focused on medical datasets, and only evaluate missingness under MCAR (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008; Shahbazian and Greco, 2023). Further, most proposed methods only evaluate imputation quality, using metrics such as MAE, MSE, RMSE, and AUC (Hasan et al., 2021; Joel and Doorsamy, 2022), although some also evaluate overall predictor accuracy  (Lin and Tsai, 2020). Additionally, the performance of MVI techniques under multi-mechanism missingness (Zhou et al., 2024) and missingness shifts (Zhou et al., 2023) remains unexplored in prior work, despite these conditions being more likely to occur in practice due to distribution shifts in production deployments (Graham, 2009).
Notably, overwhelming evidence in the literature indicates that there is no single “best-performing” MVI approach on accuracy  (Hasan et al., 2021; Lin and Tsai, 2020; Shadbahr et al., 2023; García Laencina et al., 2010; Farhangfar et al., 2008), and that model performance (narrowly measured based on ‘correctness’ thus far) depends on dataset characteristics such as size and correlation between variables (Alabadla et al., 2022) and missingness rates in the train and test sets (Lin and Tsai, 2020; Shadbahr et al., 2023).
@@
There has been some recent interest in studying the social harm that can come from poorly chosen MVI techniques (Caton et al., 2022; Zhang and Long, 2022, 2021a, 2021b; Jeong et al., 2022a; Wang and Singh, 2021; Martínez-Plumed et al., 2019; Feng et al., 2024). Most empirical studies (Wang and Singh, 2021; Zhang and Long, 2021b, 2022; Caton et al., 2022; Jeong et al., 2022b) have worked with the COMPAS (Larson et al., 2016) and Adult (Dua and Graff, 2019) datasets, the latter of which has been “retired” from community use due to issues with provenance (Ding et al., 2021). Further, these experimental studies employ randomly-generated missingness: usually by randomly sampling or using a fixed set of columns, and randomly picking rows in which to replace values with null. We critique this approach, since detecting and mitigating unfairness requires broader socio-technical thinking, such as having higher rates of missingness for minority groups and in features that are highly correlated with sensitive attributes (called proxy variables in the fairness literature) (Caton et al., 2022).
@@
A notable exception is Martínez-Plumed et al. (2019), who map social mechanisms such as prejudicial access and self-reporting bias to missingness categories like missing-by-design and item non-response. They also analyze feature correlations to study the effects of different missingness types. We adopt a similar methodology to simulate realistic missingness in this work but identify conceptual limitations in their fairness framing. The authors state: “The surprising result was to find that, […] the examples with missing values seem to be fairer than the rest.” However, asserting that some rows of data are more or less fair is misguided, as fairness is not a property of individual samples (e.g., job applicants) but of the model (e.g., in hiring decisions), which determines fairness through inclusion or exclusion in positive outcomes. We reinterpret their findings to suggest that excluding samples with missing values can increase model unfairness, reinforcing the case against deletion as a missing data strategy.
@@
Zhang and Long (2021b) evaluate MVI methods on imputation fairness, defined as the difference in imputation accuracy between privileged and disadvantaged groups. They find that imputation unfairness increases with higher missingness disparity, higher overall missingness rates, and greater data imbalance across groups. Further, they find that varying missingness mechanisms for the same imputation method impacts prediction fairness. Their analysis is limited to randomly generated null values in COMPAS. We extend their work to additional datasets, missingness scenarios, and alternative imputation fairness definitions.
@@
In a follow-up work, Zhang and Long (2022) introduce imputation fairness risk and provide bounds for “correctly specified” imputation methods. While this is a commendable theoretical contribution in a largely unexplored area, we question its broader implications: imputation quality metrics do not fully capture downstream model performance (Zhang and Long, 2021a). In other words, a classifier can perform well despite poor imputation quality (Shadbahr et al., 2023). This raises a key question: Does minimizing imputation unfairness reduce model unfairness? Our empirical findings suggest it does not, as discussed in Section 4.4.
Finally, Jeong et al. (2022b) propose a decision tree-based method that integrates fairness into model training while handling missing values. Their approach splits only on observed values to mitigate disparities introduced by imputation. Their evaluation is limited to MCAR scenarios (with more missingness for disadvantaged groups). In contrast, we assess more advanced MVI techniques under diverse missingness scenarios (MCAR, MAR, MNAR, and missingness shift) without applying fairness interventions. Nonetheless, we share their broader motivation of assessing and mitigating unfairness holistically throughout the data lifecycle. Future work could explore different combinations of MVI and fairness interventions.
Missingness and stability.
@@
We are not aware of any work investigating the effect of missing value imputation on model stability.